Kim and Kayne have been dating for years and are now thinking about getting married. As a financially sophisticated​ couple, they want to think through the tax implications of their potential union.

a. Suppose Kim and Kanye both earn

​$68000

​(so their combined income is

​$136000

Using the tax bracket information

calculate the combined tax bill that they would pay if they remain​ single, and compare that to the taxes they would pay if they were married and filed a joint return.

b. Now suppose that Kim and Kanye both earn

​$103000

​(so their combined income is

​$206000

Calculate the combined tax bill that they would pay if they remain​ single, and compare that to the taxes that they would pay if they were married and filed a joint return.

Data Table TABLE 1.2: Federal Income Tax Rates and Brackets for Individual and Joint Returns (Due by April 15, 2015) Taxable Income Tax Rates 10.0% 15.0% 25.0% 28.0% 33.0% 35 .0% 39.6% Individual Returns $0 to $9,075 $9,076 to $36,900 $36,901 to $89,350 $89,351 to $186,350 $186,351 to $405,100 $405,101 to $406,750 Over $406,750 Joint Returns $0 to $18,150 $18,151 to $73,800 $73,801 to $148,850 $148,851 to $226,850 $226,851 to $405,100 $405,101 to $457,600 Over $457,600
Answer #1

(a) Individual Incomes = $68000, Total Income = $136000

If they pay taxes individually,
Tax liability for each = 9075 * 10% + (36900 - 9075) * 15% + (68000 - 36900) * 25% = $12856.25
Total Paid by them = 12856.25 * 2 = $25712.5

If they pay joint taxes,
Total Tax liability = 18150 * 10% + (73800 - 18150) * 15% + (136000 - 73800) * 25% = $25712.5

(b) Individual Incomes = $103000, Total Income = $206000

If they pay taxes individually,
Tax liability for each = 9075 * 10% + (36900 - 9075) * 15% + (89350 - 36900) * 25% + (103000 - 89350) * 28% = $22015.75
Total Paid by them = 22015.75 * 2 = $44031.5

If they pay joint taxes,
Total Tax liability = 18150 * 10% + (73800 - 18150) * 15% + (148850 - 73800) * 25% + (206000 - 148850) * 28% = $44927